# Paint Outside the Box: Synthesizing and Selecting Training Data for Visual Grounding

### v1

@@ Abstract
Visual grounding aims to localize the image regions based on a textual query. Given the difficulty of large-scale data curation, we investigate how to effectively learn visual grounding under data-scarce settings in this paper. To address data scarcity, we propose a novel framework, POBF (Paint Outside the Box, then Filter). POBF synthesizes images by inpainting outside the box, tackling a label misalignment issue encountered in previous works. Furthermore, POBF leverages an innovative filtering scheme to identify the most effective training data. This scheme combines a hardness score and an overfitting score, balanced by a penalty term. Experimental results show that POBF achieves superior performance across four datasets, delivering an average improvement of 5.83% and outperforming leading baselines by 2.29% to 3.85% in accuracy. Additionally, we validate the robustness and generalizability of POBF across various generative models, data ratios, and model architectures.
1 Introduction
Visual grounding [39, 41, 93, 11, 29, 90, 63, 12] is a fundamental task in vision-language understanding. The objective is to locate a specific image region based on a free-form textual query. However, constructing visual grounding datasets [92, 47, 50, 30] with dense region-text annotations is both time-intensive and costly, which hinders the progress in this field. Motivated by the difficulty of large-scale data curation, we investigate the feasibility of effectively learning visual grounding under data-scarce settings, such as with access to only 1% training data.
With recent advances in image generation and captioning models [36, 54, 55, 40], a natural approach to address data scarcity is to generate training data. The advanced generative models, extensively pretrained on large-scale real-world data, have captured the correspondence between natural images and their textual descriptions. It is probable to benefit visual grounding by distilling the knowledge from generative models through training data generation. Recent works have explored the generation of densely annotated data to train models for fine-grained understanding tasks, such as object detection and image segmentation [96, 14, 83, 16, 38, 15, 7, 4]. However, these methods rely on specialized networks trained on extensive dense annotations, which hinders their flexibility and generalizability.
To address this limitation, we propose leveraging off-the-shelf generative models, trained solely on image-caption pairs in a weakly-supervised manner, to synthesize training data. Our method offers two advantages. First, the use of the off-the-shelf generative models allows our method to generalize across a wider range of domains, since the generative models are trained on web-scale image-caption data. Second, we can benefit from the rapid progress in generative models, incorporating the latest advancements into our data generation pipeline without additional training on dense annotations.
@@ 1 Introduction
Figure 1: Examples of generated images from different methods. X-Paste [96] synthesizes unnatural images. Gen2Det [64] exhibits the label misalignment issue. GeoDiffusion [7] either fails to generate the correct object or introduces excessive artifacts. In contrast, our method synthesizes high-quality images with accurately aligned bounding boxes.
@@
Without being trained on dense annotations, the off-the-shelf generative models struggle to directly generate training data with dense annotations. To address this, we utilize the region-text annotations from limited real data to generate training data through inpainting. However, naively inpainting the object within the bounding box [64] can lead to label misalignment issues. Specifically, the generated object may not spatially align well with the box or even fail to appear entirely. Figure 1 shows the examples. The synthetic ‘right cow’ in 1(a) misaligns with the box, and the ‘batter’ in 1(b) disappears. To mitigate this problem, we propose a strategy that preserves the content within the box while generating the background outside. Figure 1 shows we can generate high-quality synthetic images that align well with the original annotations.
@@
Moreover, training neural networks with synthetic data presents another challenge: not all samples contribute positively to model performance [4, 10, 46, 45]. To identify the most effective training data, we propose a filtering scheme consisting of three components: a hardness score, an overfitting score, and a penalty term. The hardness score prioritizes the easiest samples, as it is challenging to model outliers when the basics are not adequately captured in our data-scarce regime [62]. The overfitting score removes samples exacerbating overfitting outside the box, as gigantic models are prone to overfitting on unintended features (known as shortcut learning[18]), especially with limited training data. Furthermore, to manage the trade-off between two scores, we introduce a penalty term. Finally, we combine the selected synthetic data with real data as the training data to train the model.
We propose a novel framework named POBF (Paint Outside the Box, then Filter). This approach leverages generative models to inpaint and caption images, thereby expanding the limited training set effectively. Additionally, it incorporates a filtering scheme to select the most effective synthetic data for training. Our contributions are as follows:
@@
  • We address a challenging data-scarce scenario in visual grounding by proposing POBF, an innovative framework that synthesizes images by inpainting outside the box. POBF uses off-the-shelf data generators pretrained solely on image-caption pairs, which makes it more flexible and adaptable.
@@
  • We propose a novel filtering scheme tailored to our data-scarce setting to identify the most effective synthetic training samples. This scheme comprises a hardness score, an overfitting score, and a penalty term.
  • Experimental results demonstrate that our proposed framework yields a 5.83% improvement across four datasets. Additionally, POBF surpasses leading baselines, such as X-Paste [96] and GeoDiffusion [7], with an accuracy increase ranging from 2.29% to 3.85%. We validate the robustness and generalizability of POBF across generative models, data ratios, and model architectures.
2 Problem Definition
Let $I$ and $T$ denote an input image and a natural language description, respectively. Visual grounding is the task of identifying the target region $R\subseteq I$ within the image that the description refer to. Typically, $R$ is represented as a bounding box $B=(x,y,w,h)$, where $(x,y)$ specifies the box center coordinates, and $w$ and $h$ define its width and height. In this paper, we formulate the problem of learning visual grounding from data-scarce scenarios. Given a dataset $D$, we aim to train a model effectively using only a small subset comprising $N$of the entire data, where $N\%\ll 1$.
3 Methodology
@@
Figure 2: Overview of our proposed framework POBF, which consists of four key steps: data generation, teacher training, data filtering, and student training. Here, $(I,T,B)$ represents a real data sample, while $T^{\prime}$ and $I^{\prime}_{j}$ denote the generated caption and image, respectively. $\mathcal{G}_{T}$ and $\mathcal{G}_{I}$ are the generative models. $\mathcal{T}$ and $\mathcal{S}$ refer to the teacher and student model, repectively.
@@
In this paper, we introduce a novel, model-agnostic framework called POBF (Paint Outside the Box, then Filter). An overview of our pipeline is provided in Figure 2. Our approach begins with a data generation process, where we synthesize images and captions using image inpainting and captioning techniques. This process utilizes the image generator and captioner pretrained on large-scale image-caption pairs. Next, we train a teacher model using the limited real data. We then apply a filtering scheme to select effective synthetic images. Specifically, the well-trained teacher network assigns a hardness score, an overfitting score, and a penalty term to each synthetic image. Finally, we retain the synthetic images with high scores and combine them with real data to train the student model. The proposed POBF framework is simple, effective, and does not rely on any auxiliary models pretrained with dense annotated data.
3.1 Cross-modality Generation
@@
Image Generation with Inpainting. We begin by synthesizing images through image inpainting, a process that reconstructs masked regions within an image based on a given prompt, while preserving the rest of the image. To address the label misalignment problem, we propose inpainting outside of the bounding box instead. For each data point ($I,T,B$), we first use a captioning model $\mathcal{G}_{T}$ to generate a caption $C$ for $I$. The bounding box $B$, image $I$, and caption $C$ are then input into the inpainting model $\mathcal{G}_{I}$, yielding an generated image $I^{\prime}$. As shown in Figure 2, the red bus within the bounding box remains unchanged, allowing it to continue serving as the ground truth. The synthetic image $I^{\prime}$ can be paired with the original text $T$ and bounding box $B$ to form a new training sample ($I^{\prime},T,B$). By generating $K$ synthetic images per real sample, denoted as $\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{K}\}$, the training dataset effectively increases in size by a factor of $K$.
We utilize Stable Diffusion XL [54] as the inpainting model, which is pretrained on image-caption pairs with the objective of predicting masked regions. To increase diversity, we set a large strength parameter, which controls the amount of noise added to the base image thereby affects the similarity of the output to the base image.
@@
Text Generation with Captioning. To augment the original text $T$, we feed the cropped image within the bounding box $B$ to the captioning model $\mathcal{G}_{\text{T}}$, which generates a new caption $T^{\prime}$ describing the object. The generated caption $T^{\prime}$ can be paired with $B$ and either the original image $I$ or the generated image $I^{\prime}$ to create a new training sample. In this work, we employ BLIP [36] as the image captioner, with nucleus sampling [24] to improve diversity.
3.2 Filtering Scheme
@@
Recognizing that not all synthetic samples contribute positively to model performance, we propose a novel filtering scheme to select the most effective images for training. Initially, we use the accessible real training dataset to train a teacher model $\mathcal{T}$. Then, $\mathcal{T}$ evaluates synthetic images based on two scores and one penalty term to determine their effectiveness as training samples. We provide an illustration in Figure 3 to demonstrate the input used when calculating each score for a generated image. For clarity in the following discussions, we assume a real data sample $(I,T,B)$ and denote the set of images generated from it as $\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{K}\}$.
@@
Hardness Score. We propose to filter harmful synthetic images via computing hardness. Intuitively, in our data-scarce regime, it is difficult to model outliers effectively as fundamental patterns may not be fully captured [62]. Since the model require easiest examples to provide coarse-grained information about the target function, we prioritizes synthetic images with lower losses. We define the hardness score for each of the synthetic image $I_{j}^{\prime}$, where $j\in\{1,2,\dots,K\}$, as follows:
$S_{1}(I_{j}^{\prime})=\text{IOU}(\mathcal{T}(I_{j}^{\prime},T),B)$ (1)
@@
Overfitting Score. Additionally, we introduce a novel score to filter out synthetic images that may hinder training by exacerbating overfitting. Given the gigantic model architecture relative to the limited number of real data, the model is prone to overfit by relying on image background features rather than the actual object in the bounding box. To mitigate this issue, we assess the potential of each generated image to contribute to overfitting. Specifically, we modify each generated image by masking the area within the bounding box and input this variant into $\mathcal{T}$. If $\mathcal{T}$ can still correctly predict the bounding box, it indicates that $\mathcal{T}$ may leverage unintended features present in the generated background. Incorporating such data into training dataset risks the model learning these undesirable features, thereby exacerbating overfitting. To prevent this, we define an overfitting score to filter these problematic samples. The score for each generate image $I_{j}^{\prime}$, where $j\in\{1,2,\dots,K\}$, is defined as:
$S_{2}(I_{j}^{\prime})=1-\text{IOU}(\mathcal{T}(\text{Mask}(I_{j}^{\prime},B),T),B)$ (2)
where $\text{Mask}(I_{j}^{\prime},B)$ is a operation that sets all pixel values within the bounding box $B_{i}$ to zero.
Figure 3: An illustration of the process of calculation each score for a generated image.
@@
$P(I_{j}^{\prime})=\text{IOU}(\mathcal{T}(I_{j}^{\prime},\emptyset),B)$ (3)
where $\emptyset$ denote an empty textual string.
The final score for a generated image $I_{j}^{\prime}$ is calculated as:
$\text{Score}(I_{j}^{\prime})=\lambda_{1}S_{1}(I_{j}^{\prime})+\lambda_{2}S_{2}(I_{j}^{\prime})+\lambda_{P}P(I_{j}^{\prime})$ (4)
@@
Out of the $K$ generated images $\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{K}\}$ from each real data point ($I,T,B$), we retain one synthetic image $I^{\prime}_{m}$ with the highest score, where $m=\arg\max_{j}\text{Score}(I^{\prime}_{j})$ and $j\in\{1,2,\dots,K\}$. This yields a synthetic data point ($I^{\prime},T,B$). By incorporating the selected synthetic data points into the training set, we effectively double its size. Finally, we then train a student model $\mathcal{S}$. During training, the real text $T$ is randomly replaced with the corresponding generated caption $T^{\prime}$ with probability $q$.
4 Experiments
Dataset. RefCOCO [92], RefCOCO+ [92], and RefCOCOg [47, 50] datasets are derived from MSCOCO 2014. Following the common splits used in previous studies [11, 12], we report performance on the testA (containing multiple people) and the testB (containing multiple instances of other objects) for RefCOCO and RefCOCO+, and the umd-test splits for RefCOCOg [50]. The ReferItGame dataset [30] is collected through an online two-player game. We report performance on the testset of ReferItGame. The statistics for these four datasets are presented in Table 1. For each dataset, we randomly sample a subset comprising 1% of the entire dataset to simulate data-scarce scenarios.
Table 1: Statistics of visual grounding datasets.

### v2

@@ Abstract.
Visual grounding aims to localize the image regions based on a textual query. Given the difficulty of large-scale data curation, we investigate how to effectively learn visual grounding under data-scarce settings in this paper. To address the data scarcity, we propose a novel framework, POBF (Paint Outside the Box and Filter). POBF synthesizes images by inpainting outside the box, tackling a label misalignment issue encountered in previous works. Furthermore, POBF leverages an innovative filtering scheme to select the most effective training data. This scheme combines a hardness score and an overfitting score, balanced by a penalty term. Extensive experiments across four benchmark datasets demonstrate that POBF consistently improves performance, achieving an average gain of 5.83% over the real-data-only method and outperforming leading baselines by 2.29%–3.85% in accuracy. Additionally, we validate the robustness and generalizability of POBF across various generative models, training data sizes, and model architectures.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; isbn: ;
@@
Figure 1. Examples of generated images from different methods. X-Paste (Zhao et al., 2023) synthesizes unnatural images. Gen2Det (Suri et al., 2024) exhibits the label misalignment issue. GeoDiffusion (Chen et al., 2024d) either fails to generate the correct object or introduces excessive artifacts. In contrast, our method synthesizes high-quality images with accurately aligned bounding boxes.
1. Introduction
Visual grounding (Liu et al., 2019b, a; Deng et al., 2021; Kamath et al., 2021; Deng et al., 2023; Liang et al., 2023; Lin et al., 2024; Xiao et al., 2024a; Ye et al., 2023; Yu and Li, 2024; Su et al., 2024) is a fundamental task in vision-language understanding. The objective is to locate a specific image region based on a free-form textual query. However, constructing visual grounding datasets with dense region-text annotations is both time-intensive and costly, which hinders the progress in this field. Motivated by the difficulty of large-scale data curation, we investigate the feasibility of effectively learning visual grounding under data-scarce settings, where only few real training data are accessible.
With recent advances in image generation and captioning (Li et al., 2022; Podell et al., 2023; Rombach et al., 2022; Liu et al., 2024), a natural approach to address data scarcity is to synthesize training data. These powerful generative models, pretrained on large-scale real-world data, have captured the correspondence between images and textual descriptions. It is probable that distilling such knowledge can benefit visual grounding through training data generation. Many recent works (Zhao et al., 2023; Fan et al., 2024; Yang et al., 2024; Feng et al., 2024; Li et al., 2023; Fang et al., 2024; Chen et al., 2024d, e) have explored the generation of densely annotated data to train models for fine-grained understanding tasks, such as object detection and image segmentation. However, they often rely on specialized networks trained on extensive dense annotations, which limits their flexibility and generalizability.
To overcome this limitation, we propose to leverage off-the-shelf generative models, trained solely on image-caption pairs in a weakly-supervised manner, for training data generation. Compared to the methods relying on specialized networks, our method offers two advantages. First, our method could generalize across a wider range of domains, since the generative models we used are trained on web-scale image-caption data. Second, we can benefit from the rapid progress in generative models, incorporating the latest advancements into our data generation pipeline without additional training on dense annotations.
@@ 1. Introduction
Although the off-the-shelf generative models are capable of synthesizing images across diverse domains, they struggle to directly generate data with region-text labels, as they are not trained with dense annotations. Therefore, we leverage the limited region-text annotations available in real data to generate training data through inpainting. However, a naive approach—inpainting the object within the bounding box (e.g., Gen2Det (Suri et al., 2024), see Figure 1)—can result in label misalignment issues. Specifically, the generated object may not spatially align well with the box or even fail to appear entirely. For example, the synthetic ‘right cow’ in Figure 1(a) misaligns with the box, and the ‘batter’ in Figure 1(b) disappears. To tackle the label misalignment problem, we propose a paint-outside-the-box strategy, which preserves the content within the box while generating the background outside. This results in higher-quality images that align well with the original annotations.
@@
Moreover, training neural networks with synthetic data presents another challenge: not all samples contribute positively to the model performance (Chen et al., 2024e, a; Malach and Shalev-Shwartz, 2017; Mahmoud et al., 2024). To identify the most effective training data, we propose a filtering scheme consisting of three components: a hardness score, an overfitting score, and a penalty term. The hardness score prioritizes the easiest samples, as it is challenging to model outliers when the basic patterns are not adequately captured in our data-scarce regime (Sorscher et al., 2022). The overfitting score penalizes samples exacerbating overfitting the unintended features outside the box, as shortcut features from the background could lead the gigantic model to overfit (Geirhos et al., 2020; Beery et al., 2018; Xiao et al., 2020), especially with limited training data. Furthermore, to manage the trade-off between two scores, we introduce a penalty term. Finally, we combine the selected synthetic data with the real data to train our visual grounding model.
In summary, we propose a novel framework named POBF (Paint Outside the Box and Filter). POBF synthesizes high-quality training samples by image inpainting and captioning, then employs an effective filtering scheme for data selection. Our contributions are:
@@
  • We address a challenging problem of visual grounding in data-scarce scenarios by introducing POBF, a flexible and adaptable framework that leverages off-the-shelf data generators pretrained solely on image–caption pairs.
  • We propose a novel data generation strategy for visual grounding, paint outside the box, which overcomes the label misalignment issues in previous approaches.
@@
  • We introduce a new filtering scheme designed for data-scarce settings to identify the most effective synthetic training samples. This scheme integrates a hardness score, an overfitting score, and a penalty term.
  • Experiments demonstrate that POBF improves performance by 5.83% over real-data-only training and outperforms strong baselines by a 2.29%–3.85% margin. We further validate its robustness and generalizability across generative models, training data sizes, and model architectures.
2. Related Work
Visual Grounding. Early approaches addressed visual grounding in a two-stage framework (Liu et al., 2019b; Yu et al., 2018; Liu et al., 2019a), which heavily depended on the quality of proposals and large computational costs due to the large number of proposals. To overcome the limitations, more efficient one-stage methods have been developed (Deng et al., 2021; Kamath et al., 2021; Ye et al., 2022; Su et al., 2023; Deng et al., 2023; Xiao et al., 2023; Shi et al., 2023; Xiao et al., 2024b; Yang et al., 2022b; Yao et al., 2024; Chen and Li, 2022; Jiao et al., 2023; Weerakoon et al., 2022). TransVG (Deng et al., 2021) incorporates a DETR encoder for visual feature extraction and introduces the first transformer-based framework for visual grounding. QRNet (Ye et al., 2022), VG-LAW (Su et al., 2023), and TransVG++ (Deng et al., 2023), integrate language-guided knowledge within the intermediate layers of visual backbones to compute query-dependent visual attention dynamically. As the first diffusion-based framework, LG-DVG (Chen and Li, 2023) progressively reasons the queried object by denoising noisy boxes with the language guide.
With the development of pretrained large multimodal models, several studies (He et al., 2024a; Yang et al., 2023a; Han et al., 2024; Kang et al., 2025; Yan et al., 2024) have leveraged these foundation models to localize objects. For example, (He et al., 2024a; Yang et al., 2023a) employ GradCAM (Selvaraju et al., 2017) to produce gradient-based visual explanations, represented as heatmaps for grounding outputs. (Kang et al., 2025) demonstrates that only a few attention heads in large vision-language models can provide accurate and interpretable text-image attention maps for object localization. Inspired by reinforcement learning, ViGoR (Yan et al., 2024) integrates a reward model for visual grounding outputs, then use the reward model for instruction tuning pretrained multimodal models. In parallel, other works (Chen et al., 2024c, 2023; Zhao et al., 2024; Zeng et al., 2024) focus on training multimodal large language models for visual grounding using extensive, fine-grained labeled datasets. Furthermore, recent research has expanded traditional visual grounding to gigapixel-level (Ma et al., 2024) and sensing (Sun et al., 2022) domain images.
Previous works typically train models on large-scale labeled visual grounding datasets. To mitigate the issue of data scarcity, several studies have explored unsupervised (Wang et al., 2024c), weakly-supervised (Chen et al., 2024b; Liu et al., 2021; Sun et al., 2021; Jin et al., 2023; Liu et al., 2023), and semi-supervised (Sun et al., 2023) learning methods. However, these methods still rely heavily on the availability of large amount of in-domain real data. Instead, we explore a more constrained data-scarce setting, where only a limited amount of real visual grounding data exists. We further propose a data generation strategy and filtering scheme for data-scarce visual grounding.
@@ 2. Related Work
Figure 2. Overview of our proposed framework POBF, which consists of four steps: data generation, teacher training, data filtering, and student training. The real data sample (image, caption, box) is denoted as $(I,T,B)$. $T^{\prime}$ and $I^{\prime}_{j}$ denote the generated caption and image, respectively. $\mathcal{G}_{T}$ and $\mathcal{G}_{I}$ are generative models. $\mathcal{T}$ and $\mathcal{S}$ refer to the teacher and student model, respectively.
Training with Synthetic Data. Building off image generation and image captioning, training on synthetic data has emerged as a promising research direction (Yang et al., 2023b; Rotstein et al., 2024; Xuan et al., 2023; Tian et al., 2023; Du et al., 2023). In fine-grained understanding tasks, researchers explore the use of synthetic data in object detection (Fang et al., 2024; Wang et al., 2024a; Chen et al., 2024d; Zhang et al., 2023; Feng et al., 2024; Suri et al., 2024), image segmentation (Yang et al., 2024; Li et al., 2023; Xie et al., 2024; Fan et al., 2024; Zhao et al., 2023), and visual grounding (He et al., 2024b; Wang et al., 2024b; Jiang et al., 2022). Some methods (Li et al., 2023; Zhang et al., 2023; Feng et al., 2024) incorporate task-specific heads into image generators and fine-tune them with extensive data to synthesize images paired with annotations. Others (Zhao et al., 2023; Fan et al., 2024; He et al., 2024b) first generate images, then employ an auxiliary model to annotate the synthetic data. For instance, DiverGen (Fan et al., 2024) generates images by a vanilla diffusion model and subsequently uses SAM (Kirillov et al., 2023) for dense annotation. Several approaches (Fang et al., 2024; Chen et al., 2024d; Yang et al., 2024; Wang et al., 2024a) employ visual priors (e.g. segmentation masks or edges) to guide the synthetic image generation process via controllable diffusion models. For example, FreeMask (Yang et al., 2024) uses FreestyleNet (Xue et al., 2023) to synthesize images conditioned on target dataset-provided masks. However, these methods rely on either the data generator, the annotator, or both to be pretrained on large-scale densely annotated datasets.
Some prior works (Jiang et al., 2022; Wang et al., 2024b; He et al., 2024b; Zheng et al., 2024) have focused on generating labeled training data for visual grounding, but they typically rely on specialized networks—such as object detectors or image generators trained on extensively annotated datasets—which limits their generalizability and flexibility. Departing from these approaches, we introduce a novel training data generation method that leverages off-the-shelf generative models trained solely on image–caption pairs, eliminating the reliance on extensive dense annotations and making our approach more flexible and scalable. In addition, we incorporate a filtering scheme tailored for inpainted images.
Data Pruning. Recent advancements in data selection methods have gained significant attention from the research community (Xie et al., 2023; Jiang et al., 2024; Mahmoud et al., 2024). These methods typically employ pre-defined criteria to assign a scalar score to each training example. While simple heuristics like repetition (Lee et al., 2021), string existence (Penedo et al., 2023), or semantic similarity (Abbas et al., 2023) are commonly used, more sophisticated criteria have also been explored, including gradient norms (Paul et al., 2021), forgetfulness (Toneva et al., 2018), and influence function scores (Yang et al., 2022a). A common strategy is to prioritize data samples with lower training loss (Yao et al., 2020; Chen et al., 2019). This approach leverages the observation that DNNs initially learn generalizable patterns before gradually overfitting to noisy data (Song et al., 2019; Li et al., 2020b). To counter confirmation bias, techniques such as co-training models (Yu et al., 2019; Li et al., 2020a) have been proposed, where two networks are trained concurrently and use their combined predictions for selection. MentorNet (Jiang et al., 2018) further extends this idea by introducing a learned curriculum to guide the selection of training data.
Our use of separately trained teacher networks contrasts with recent studies (Chen et al., 2024e; Rose et al., 2023; Yang et al., 2022c; Tang et al., 2023; He et al., 2024b) on synthetic training data selection, which filter out text-image pairs with low CLIP similarity scores. In contrast to prior work, we propose to identify the most effective synthetic training data from the hardness and overfitting perspectives, specifically tailored to the problem of data-scarce visual grounding.
3. Problem Definition
Let $I$ and $T$ denote an input image and a natural language description, respectively. Visual grounding is the task of identifying the target region $R\subseteq I$ within the image that $T$ refers to. Typically, $R$ is represented as a bounding box $B=(x,y,w,h)$, where $(x,y)$ specifies the box center coordinates, and $w$ and $h$ define its width and height. In this paper, we investigate the problem of learning visual grounding in data-scarce scenarios, namely data-scarce visual grounding, where only a very limited amount of annotated real training data is available—e.g., as little as 1% of the dataset.
4. Methodology
@@
To address the data-scarce visual grounding problem, we introduce a novel, model-agnostic framework called POBF (Paint Outside the Box and Filter). As illustrated in Figure 2, POBF consists of four stages. (a) Our approach begins with a data generation process, where we synthesize images and captions using image inpainting and captioning techniques. This process utilizes the image generator and captioner pretrained on large-scale image-caption pairs. (b) Next, we train a teacher model using the limited real data. (c) We then apply a filtering scheme to select effective synthetic images. Specifically, the well-trained teacher network assigns a hardness score, an overfitting score, and a penalty term to each synthetic image. (d) Finally, we retain the synthetic images with high scores and combine them with real data to train the student model. The proposed POBF framework is simple, effective, and does not rely on any auxiliary models pretrained with dense annotated data.
4.1. Cross-modality Data Generation
@@
Image Generation with Inpainting. We begin by synthesizing images through image inpainting, a process that regenerate masked regions within an image based on a given prompt, while keeping the unmasked areas intact. To address the label misalignment problem, we propose to paint outside of the bounding box instead. For each data point ($I,T,B$), we first use a captioning model $\mathcal{G}_{T}$ to generate a caption $C$ for $I$. The bounding box $B$, image $I$, and caption $C$ are then input into the inpainting model $\mathcal{G}_{I}$, yielding an generated image $I^{\prime}$. As shown in Figure 2(a), the red bus within the bounding box remains unchanged, making it strictly align with the ground truth bounding box after generation. The synthetic image $I^{\prime}$ can be paired with the original text $T$ and bounding box $B$ to form a new training sample ($I^{\prime},T,B$). By generating $K$ synthetic images per real sample, denoted as $\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{K}\}$, the training dataset effectively increases in size by a factor of $K$. A key hyperparameter in the inpainting process is strength, which controls the amount of noise added to the base image and, consequently influences the similarity between the generated output and the original image. To encourage data diversity, we set the strength parameter to a high value.
@@
Text Generation with Captioning. To augment the original text $T$, we feed the cropped image within the bounding box $B$ to the captioning model $\mathcal{G}_{\text{T}}$, which generates a new caption $T^{\prime}$ describing the object. The generated caption $T^{\prime}$ can be paired with $B$ and either the real image $I$ or the generated image $I^{\prime}$ to create a new training sample. In this work, we employ $p$-sampling to improve diversity.
4.2. Filtering Scheme
Figure 3. Illustration of the inputs used to compute each score for a generated image. Here, $B$ and $\hat{B}$ denote the ground truth and predicted bounding box, respectively.
@@ 4.2. Filtering Scheme
Recognizing that not all synthetic samples contribute positively to model performance (Chen et al., 2024e, a; Malach and Shalev-Shwartz, 2017; Mahmoud et al., 2024), we propose a novel filtering scheme to select the most effective images for training. Initially, we use the limited real training dataset to train a teacher model $\mathcal{T}$. Then, $\mathcal{T}$ evaluates synthetic images based on two scores and one penalty term to determine their effectiveness as training samples. We provide an illustration in Figure 3 to demonstrate the input used when calculating each score for a synthetic image. For clarity in the following discussions, we assume a real data sample $(I,T,B)$ and denote the set of images generated from it as $\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{K}\}$.
@@
Hardness Score. We propose to filter harmful synthetic images via computing hardness. Intuitively, in our data-scarce regime, it is difficult to model outliers effectively as fundamental patterns may not be fully captured (Sorscher et al., 2022). Since the model require easiest examples to provide coarse-grained information about the target function, we prioritize synthetic images that are less challenging for the teacher model $\mathcal{T}$. We define the hardness score for each of the synthetic image $I_{j}^{\prime}$, where $j\in\{1,2,\dots,K\}$, as follows:
(1) $S_{1}(I_{j}^{\prime})=\text{IOU}(\mathcal{T}(I_{j}^{\prime},T),B)$
@@
Overfitting Score. Additionally, we introduce another score to filter out synthetic images that may hinder training by exacerbating overfitting. Given the gigantic model architecture relative to the limited number of real data, the model is prone to overfit by relying on image background features rather than the actual object in the bounding box (Beery et al., 2018; Xiao et al., 2020). To mitigate this, we assess the potential of each generated image to contribute to overfitting. Specifically, we modify each generated image by masking the area within the bounding box and input this variant into $\mathcal{T}$. If $\mathcal{T}$ can still correctly predict the bounding box, it indicates that $\mathcal{T}$ may leverage unintended features present in the generated background. Incorporating such data into training dataset risks the model learning these unintended features, thereby exacerbating overfitting. We define an overfitting score to filter these problematic samples. The score for each generate image $I_{j}^{\prime}$, where $j\in\{1,2,\dots,K\}$, is defined as:
(2) $S_{2}(I_{j}^{\prime})=1-\text{IOU}(\mathcal{T}(\text{Mask}(I_{j}^{\prime},B),T),B)$
where $\text{Mask}(I_{j}^{\prime},B)$ is a operation that sets all pixel values within the bounding box $B$ to zero.
@@
(3) $P(I_{j}^{\prime})=\text{IOU}(\mathcal{T}(I_{j}^{\prime},\emptyset),B)$
where $\emptyset$ denote an empty textual string. This term can be regarded as the prior knowledge in the generated image.
The three scores are individually normalized to have zero mean and unit standard deviation. Then The final score for a generated image $I_{j}^{\prime}$ is calculated as:
(4) $\text{Score}(I_{j}^{\prime})=\lambda_{1}S_{1}(I_{j}^{\prime})+\lambda_{2}S_{2}(I_{j}^{\prime})+\lambda_{P}P(I_{j}^{\prime})$
@@
Out of the $K$ generated images $\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{K}\}$ from each real data point ($I,T,B$), we retain one synthetic image $I^{\prime}_{m}$ with the highest score, where $m=\arg\max_{j}\text{Score}(I^{\prime}_{j})$ and $j\in\{1,2,\dots,K\}$. This yields a synthetic data point ($I^{\prime},T,B$). By incorporating the selected synthetic data points into the training set, we effectively double its size. Finally, we train a student model $\mathcal{S}$, where the real text $T$ is randomly replaced with the corresponding generated caption $T^{\prime}$ with probability $q$ during training.
5. Experiments
Dataset. RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016), and RefCOCOg (Mao et al., 2016; Nagaraja et al., 2016) datasets are derived from MSCOCO 2014. Following the common splits used in previous studies (Deng et al., 2021, 2023), we report performance on the testA (containing multiple people) and the testB (containing multiple instances of other objects) for RefCOCO and RefCOCO+, and the umd-test splits for RefCOCOg (Nagaraja et al., 2016). The ReferItGame dataset (Kazemzadeh et al., 2014) is collected through an online two-player game. We report performance on the test set of ReferItGame. The statistics for these four datasets are presented in Table 1. For each dataset, we randomly sample subsets comprising 1% of the entire dataset to simulate data-scarce scenarios.
Table 1. Statistics of visual grounding datasets.